OpenStudy (anonymous):

Factor completely: 3a4y3 − 12a3y2 + 6a2y?

jhonyy9 (jhonyy9):

3a^2y(a^2y^2 - 4ay + 2) = 3a^4y^3 -6a^3y^2 -6a^3y^2 +6a^2y = 3a^3y^2 (ay - 2) -6a^2y (ay -1) = =3a^2y(ay(ay-2) -2(ay-1))